OpenStudy (anonymous):

Castel and gabriella are selling pies for a school fundraiser.Customers can buy apple pies and lemon meringue pies.Castel sold 6 apple pies and 4 lemon neringue pies.Gabriella sold 6 apple pies and 5 lemon meringue pies.What is the total cost for each of one apple pie and one lemon meringue pies can someone help me solving this by using the elimination method?

OpenStudy (anonymous):

so ok you mean like... 6a+4a=x 6a+5l=x ?

OpenStudy (anonymous):

and then elimination method? to that system of equations?

OpenStudy (anonymous):

ok looking at this though is there a total cost to how much EACH person made? this problem looks incomplete

OpenStudy (anonymous):

yes do you know how to solve it......HOW?

OpenStudy (anonymous):

I do this problem looks incomplete.

OpenStudy (anonymous):

like how can we find the total cost of each product when we don't know how much they made?

OpenStudy (anonymous):

Is there more to the problem you forgot to add on?

OpenStudy (anonymous):

No ....... you have to find the total cost did you read the problem cuz the numbers are there

OpenStudy (anonymous):

Castel sold 6 apple pies and 4 lemon neringue pies.Gabriella sold 6 apple pies and 5 lemon meringue pies

OpenStudy (anonymous):

those are the only numbers you provided tho.

OpenStudy (anonymous):

yes r u a freshman

OpenStudy (anonymous):

no i just don't understand how you would find one apple pie and one lemon meringue pies if you don't know what the total outcome is

OpenStudy (anonymous):

like apple pie and lemon pie = total product but we dont have total product to work from. we just have two variables on the side when we set the problem up

OpenStudy (anonymous):

you have to create a system of equations to figure out the problem

OpenStudy (anonymous):

right but what does each one equal to? 6a+4a=x 6a+5l=x

OpenStudy (anonymous):

6a+4l=x 6a+5l=x

OpenStudy (anonymous):

you can't solve this without having the total cost spent

OpenStudy (anonymous):

yes kellie that's what i'm saying.

OpenStudy (anonymous):

@kelliegirl33

OpenStudy (anonymous):

unless the two total products are supposed to be set equal to each other?

OpenStudy (anonymous):

like: 6a+4a=6a+5l?

OpenStudy (anonymous):

i'm trying to tell y'all that you solve by eliminating the equations

OpenStudy (anonymous):

you can't use eliminating the equations without a total product from each equation http://www.mathportal.org/algebra/solving-system-of-linear-equations/elimination-method.php

OpenStudy (anonymous):

you cannot eliminate

OpenStudy (anonymous):

Is there more to this problem ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i'm sorry i really do want to help but you cant solve that problem unless you know the total spent for each person.

OpenStudy (anonymous):

you cannot find the cost of one product if you do not have any money value in the problem, such as the total cost

OpenStudy (anonymous):

um i did leave out 1 part i'm SO SORRY

OpenStudy (anonymous):

IT'S OK just give it to us so that we can help you

OpenStudy (anonymous):

CASTEL'S TOTAL COST IS 80 DOLLARS and GABRIELLA'S TOTAL COST IS 94 DOLLARS I'M SORRY I'M TIRED

OpenStudy (anonymous):

jessie....I am not gonna take over this problem....so..um...you got this ?

OpenStudy (anonymous):

yes i do :)

OpenStudy (anonymous):

it's ok tionne i'll walk you through it

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

do you need help ?

OpenStudy (anonymous):

6a+4l=80 6a+5l=94

OpenStudy (anonymous):

ok first -(6a+4a=80) 6a+5l=94 6a+5l=94 -6a-4l=-80

OpenStudy (anonymous):

so you add

OpenStudy (anonymous):

& y is it negative

OpenStudy (anonymous):

you need to reverse the equation so that the variable can cancel out. then you can find l. after finding l you can input that number into the original equation to find the value of a.

OpenStudy (anonymous):

6a + 4l = 80 6a + 5l = 94 --->(-1) ----------- 6a + 4l = 80 -6a - 5l = -94 (result of multiplying by -1) ------------add -l = -14 l = 14

OpenStudy (anonymous):

yeee and now you just have to add 14 for l into the original equation so that you can find a!

OpenStudy (anonymous):

so you add 4 +14 and 5 + 14 right

OpenStudy (anonymous):

uhm ok i mean sub 14 for l in any of the equations

OpenStudy (anonymous):

like 6a + 4(14)=80

OpenStudy (anonymous):

that will help you find a

OpenStudy (anonymous):

you multiply

OpenStudy (anonymous):

yes multiply 6a+4(14)=80 6a+56=80 6a=24 a=4

OpenStudy (anonymous):

after you got that doesn't it have to be a order pair

OpenStudy (anonymous):

nicely done jessie

OpenStudy (anonymous):

thank you @kelliegirl33 and tionne it can be but they only asked for for total cost of each one, so i mean 4$ per apple pie and 14$ per lemon pie should be an adequate enough answer

OpenStudy (anonymous):

@kelliegirl33 opinion on that?^

OpenStudy (anonymous):

she left but thanks jessie um if you don't mind if you have time to help me with one more please?

OpenStudy (anonymous):

I will most definitely take a look

OpenStudy (anonymous):

ok so The school that imani goes to is selling tickets for a annual dance competition.On the first day of the ticket sales the school sold 3 senior tickets for a total of 69 dollars.The school took in 91 dollars on the second day by selling 3 senior tickets and 3 child tickets. What is the price of one senior ticket and one child ticket and its using the elimination again

OpenStudy (anonymous):

it does not have to be an ordered pair....just post your answer like jessie did above :)

OpenStudy (anonymous):

3s = 69 (solve for s and get the cost of a senior ticket. 3s + 3c = 91 (sub in what you got for s in the above equation and then solve for c.

OpenStudy (anonymous):

3s=69 3s+3c=91 find s first: 69/3=23 plug it into the second equation and find c 3(23)+3c=91 69+3c=91 3c=22 c=22/3=7.33

OpenStudy (anonymous):

o god

OpenStudy (anonymous):

do you need anymore help?

OpenStudy (anonymous):

hold on god

OpenStudy (anonymous):

why did you divide

OpenStudy (anonymous):

what you do to one side you must do to the other. i will go step by step. sorry i did geo last year and i could leave out things like subtracting and stuff. i just had to imply that i did it. 3(23)+3c=91 69+3c=91 -69 -69 3c=22 3c/3=22/3 c=22/3 c=7.33 was that clearer?

OpenStudy (anonymous):

yes bless me

OpenStudy (anonymous):

hahaha i'm glad you understand. math takes a lot of observing. is that all though?

OpenStudy (anonymous):

yes thanks jessie your name reminds me of the show but that's it thank you & have a goodnight :)

OpenStudy (anonymous):

hahaha i get that a lot and you're welcome have a great night as well!